Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

gmat prep

Expert replies
by rommysingh » Mon Sep 14, 2015 5:33 pm
if the balls are numbered from 1-100. and three balls are to be taken out by replacement then what is the probability that their sum will be odd.

1/4
3/4
1/2
5/8
6/8
Join the discussion
Source: — Problem Solving |

by Max@Math Revolution » Mon Sep 14, 2015 10:27 pm
Forget conventional ways of solving math questions. In PS, IVY approach is the easiest and quickest way to find the answer.


if the balls are numbered from 1-100. and three balls are to be taken out by replacement then what is the probability that their sum will be odd.

A. 1/4
B. 3/4
C. 1/2
D. 5/8
E. 6/8




Since the sum is odd it should be (odd+odd+odd) or (odd+even+even). The probability to choose odd is always 1/2 since it is taken out with replacement. So the probability to choose (odd+odd+odd) is 1/2 * 1/2 * 1/2 = 1/8. Similarly the probability to choose (odd+even+even) is 1/2 * 1/2 * 1/2 = 1/8. But it could be (even+odd+even) or (even+even+odd). So the probability that their sum would be odd is 1/8 + 1/8 *3 = 4/8 = 1/2. The answer is, therefore, C

Another approach.
Since the chance to choose even or odd is equal(there are 50 odd balls and 50 even balls), the probability that their sum would be odd and the probability that their sum would be even are the same, moreover they are complementary events. So it should be 1/2 without tedious calculations.




Math Revolution : Finish GMAT Quant Section with 10 minutes to spare
The one-and-only World’s First Variable Approach for DS and IVY Approach for PS with ease, speed and accuracy.
Unlimited Access to over 120 free video lessons - try it yourself (https://www.mathrevolution.com/gmat/lesson)
See our Youtube demo (https://www.youtube.com/watch?v=R_Fki3_2vO8)
Join the discussion

by [email protected] » Tue Sep 15, 2015 8:35 am
Hi rommysingh,

These types of questions can be solved in a couple of different ways, depending on how you like to organize your information and how you "see" the question.

Since half the numbered balls are 'odd' and half are 'even', and we're replacing each ball after selecting it, on any given selection we have a 50/50 chance of getting an odd or even. This means that there are....

(2)(2)(2) = 8 possible arrangements when selecting 3 balls from the box:

EEE
EEO
EOE
OEE

OOO
OOE
OEO
EOO

Since we have the same number of Odds and Evens, each of these options has the same 1/8 probability of happening. Now you just have to find the ones that give us the ODD SUM that we're asked for. They are:

EEO
EOE
OEE
OOO

Four of the eight options. 4/8 = 1/2

Final Answer: C

GMAT assassins aren't born, they're made,
Rich
Contact Rich at [email protected]
Image
Join the discussion